# GestureHYDRA: Semantic Co-speech Gesture Synthesis via Hybrid Modality Diffusion Transformer and Cascaded-Synchronized Retrieval-Augmented Generation

### v1

@@ Abstract
While increasing attention has been paid to co-speech gesture synthesis, most previous works neglect to investigate hand gestures with explicit and essential semantics. In this paper, we study co-speech gesture generation with an emphasis on specific hand gesture activation, which can deliver more instructional information than common body movements. To achieve this, we first build a high-quality dataset of 3D human body movements including a set of semantically explicit hand gestures that are commonly used by live streamers. Then we present a hybrid-modality gesture generation system GestureHYDRA built upon a hybrid-modality diffusion transformer architecture with novelly designed motion-style injective transformer layers, which enables advanced gesture modeling ability and versatile gesture operations. To guarantee these specific hand gestures can be activated, we introduce a cascaded retrieval-augmented generation strategy built upon a semantic gesture repository annotated for each subject and an adaptive audio-gesture synchronization mechanism, which substantially improves semantic gesture activation and production efficiency. Quantitative and qualitative experiments demonstrate that our proposed approach achieves superior performance over all the counterparts. The project page can be found https://mumuwei.github.io/GestureHYDRA/.
@@ 1 Introduction
Co-speech gesture synthesis aims to produce gestures synchronized with speech audio, which enables a variety of applications in filmmaking, game design, robotics, and digital human creation [11, 48, 16, 12, 47]. Considerable enthusiasm [27, 51, 49] is devoted to exploring the elusive relationship between speech audio and holistic body movements at the early stage.
Recently, a few studies [23, 19, 42, 25, 26, 57, 9] have shifted their focus on the semantic relevance within the gestures and pursue performance improvement by involving speech textual content as an additional condition. [25, 26] leverage word embedding from text scripts to perform semantic-relevant and rhythm-relevant co-speech motion modeling. The latest studies [57, 9] prefer generative semantic retrieval via large language models (LLM) for better semantic injection. However, the semantic-gesture correlation studied in these approaches is not as clear as hand gestures with explicit instructions (e.g., numbers and directions), which can be frequently observed in our daily communication. Verifiably, hand gestures not only improve the efficiency of communication but also facilitate understanding via effective visual patterns [17, 2], which constructs a crucial component for co-speech generation. Therefore, to create life-like human behaviors that can efficiently deliver explicit information, a co-speech semantic gesture generation system is imperative.
@@ 2.1 Co-speech Gesture Generation
Human motion embodies rich and expressive information, rendering human motion analysis a longstanding and extensively studied research problem [58, so3, 7, 38]. Co-speech Gesture Generation focuses on understanding the complex relationship between speech audio and holistic body movements. Initially, researchers primarily relied on rule-based methods which required extensive manual effort to model this relationship [5, 20, 18, 41, 31]. Data-driven methods [30, 45, 53, 49, 6, 21, 59, 3, 1, 14, 36, 24] replaced rules with diverse models (such as LSTM [15], VQ-VAE [39], Diffusion [35], etc.) trained on large datasets, enabling automatic speech-gesture learning.
Some studies [51, 29, 49, 6] focus on the rhythmic consistency between speech and generated body movements, particularly gestures, while roughly controlling limited motion styles using a one-shot identity vector. Recent studies [57, 9, 26] highlighted the importance of semantic gestures in conversational scenes and collected corresponding datasets. However, the semantic gestures in these works remain coarse and ambiguous. In contrast, our proposed large-scale Streamer dataset focuses on instructional and explicit semantic gestures (e.g., numbers and directions) in real-world scenarios, filling this gap.
@@ 2.2 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) is a hybrid approach that enhances generative models with external knowledge from databases. By retrieving relevant real-time data, RAG mitigates key limitations of LLMs [10, 46], such as knowledge cut-off and hallucination. Recent works [8, 56] also apply RAG to text-to-image/video tasks by incorporating retrieved visual or textual context.
Recently, RAG has also been applied to gesture generation tasks. ReMoDiffuse [55] explores the utilize of retrieved motions to provide useful guidance for the denoising process. Co-speech generation works like Semantic Gesticulator [57] merges the retrieved gesture tokens into the initial token sequence via best merging time calculation. Similarly, SIGesture [9] inserts the retrieved gestures into the denoised latent space for better generation. Unlike these methods that directly merge or insert the retrieval results back into the original gesture sequence, our approach relies on a cascaded RAG strategy, which first retrieves a key frame from relevant segments and re-generates expected rhythm and synchronized gestures with the input audio.
@@ 3.1 Task Formulation and Preliminaries
Task Formulation. Given an audio clip $A=\{a_{i}\}^{L-1}_{i=0}$, the objective of co-speech gesture generation is to reconstruct the target human gestures $G=\{g_{i}\}^{L-1}_{i=0}$ including natural body movements and responsive hand gestures with specific semantics, where $L$ denotes the sequence length. Concretely, we follow [49] to utilize a short clip with $S$ frames $G^{s}=\{g_{i}\}^{S-1}_{i=0}$ as the initial gestures (a.k.a., seed gestures).
Preliminaries for Diffusion models. Our system is built upon MDM[37], which deploys diffusion transformer blocks for human motion modeling. The diffusion process operates as a Markov chain with $T$ steps, where noisy sequences $g_{1},...,g_{T}$ originate from real observations $g_{0}\sim q(g_{0})$. Gaussian noise is gradually applied at each step: $q\left(g_{t}\mid g_{t-1}\right):=\mathcal{N}\left(g_{t};\sqrt{1-\beta_{t}}g_{t-1},\beta_{t}\mathbf{I}\right)$. The reverse process aimed to reconstruct the original gestures from distorted gesture sequences. Unlike the image generation task [34], MDM prefers to predict the original sequence $g_{0}$ instead of the noise by using a learnable deep learning network, denoted by $\overline{g}_{0}=S_{\theta}\left(g_{t},t,c\right)$, where $c$ denotes the conditions. The training goal of this model is designed to minimize the mean squared error (MSE) between the original sequence $g_{0}$ and the predicted result $\overline{g}_{0}$, expressed as $L_{t}=\mathbb{E}_{g_{0},\epsilon,t}\left[g_{0}-S_{\theta}\left(g_{t},t,\mathrm{c}\right)\right]$.
@@ 3.2 Hybrid-Modality Diffusion Transformer
Hybrid Gesture Synthesis Training Procedure. During training, we first produce noisy gestures $G^{\prime}$ perturbed by Gaussian noise $\epsilon$ and leverage a Gesture Encoder $\bm{E}_{G}$ to extract noisy gesture latent $\mathbf{G}\in\mathbb{R}^{L\times D}$. Once the masking strategy is specified, we derive the key-frame gesture $G^{m}$ from $G$ by using the training mask $M=\{m_{i}\}^{L-1}_{i=0},m_{i}\in\{0,1\}$ and then perform channel-wise concatenation on $G^{m}$ and $M$. Then a KeyFrame Encoder $\bm{E}_{K}$ similar to $\bm{E}_{G}$ is employed to extract key-frame gesture latent, which is subsequently added to the noisy gesture latent $\mathbf{G}$ to form a fused latent $\mathbf{G^{K}}\in\mathbb{R}^{L\times D}$. In the next step, we turn our attention to the speech audio aspect. We first leverage a widely used Audio Encoder [4] $\bm{E}_{A}$ to extract audio embeddings $\mathbf{A}\in\mathbb{R}^{L\times C}$. Then we follow [6] to perform cross-modality alignment by using a simple gesture-audio fusion (GAF) module with a residual block instead of cross-attention layers [40], which consists of a layer normalization (LN) operation and a multiple-layer perception (MLP). The fused latent $\mathbf{G}^{F}\in\mathbb{R}^{L\times C}$ can be defined as:
@@
Motion-Style Injective Transformer Layer. We initialize our transformer layer following [40] and observe that generated gestures fail to preserve their original styles without further fine-tuning on individual data. Most previous approaches [49, 29] adopt one-hot embeddings for speaker identity guidance, however, they cannot provide generalized performance, especially when the number of identities keeps increasing.
Recently, a zero-shot talking face generation approach [13] built upon speech-driven facial mesh animation proposes to improve its generalization ability by using an additional motion sequence as talking style reference. We take inspiration from it and propose to involve an additional gesture sequence and newly designed injective transformer layers for a more robust motion-style injection. Specifically, the motion-style injective transformer layer shown in Fig. 2(d) adopts a standard combination of a self-attention layer and a feed-forward network (FFN) layer with two style injection layers placed right after them, respectively. The style injection layer is designed to integrate the benefits of both dynamic motion components and static motion components, where the dynamic component denotes the motion-style embedding $\mathbf{S_{d}}\in\mathbb{R}^{L\times D}$ derived from a style encoder $\bm{E}_{S}$ representing captured motion-style information from external input, while the static component is an internal learnable motion memory-bank $\mathbf{S_{s}}\in\mathbb{R}^{L_{m}\times D}$ similar as  [44, 43], memorizing all the motion-styles within the training data. The final motion-style embedding $\mathbf{S}\in\mathbb{R}^{(L_{m}+L)\times D}$ is achieved by concatenating $\mathbf{S_{d}}$ and $\mathbf{S_{s}}$ along the temporal dimension and the style-injection attention can be formulated as:
@@
Dataset System FGD $\downarrow$ $\Delta BC$downarrow SAR $\uparrow$ SMD-L1 $\downarrow$ SMD-DTW $\downarrow$
@@
Training Objectives. In addition to MSE loss $\mathcal{L}_{t}$ described in 3.1, we also incorporate the velocity loss [6] $L_{vec}$ and a 3D key point loss $L_{kp}$ to ensure similar human motion prediction to stable and accurate 3D positions, which provides effective guidance for downstream applications such as controllable video generation. Since we directly predict the original gestures $\overline{x}_{0}$ rather than the noise, it is convenient to calculate these auxiliary losses. The velocity loss and the 3D keypoint loss are both formulated using L1 distance. The velocity loss measures discrepancies in velocities, while the 3D keypoint loss quantifies the gap between predicted and ground-truth 3D keypoints. The total loss function is a weighted sum of the three losses:
@@
System FGD $\downarrow$ $\Delta BC$downarrow
@@ 3.3 Cascaded-Synchronized RAG
A recent study [57] attempts to obtain injection timestamps by detecting audio beats corresponding to the stroke phase of gestures and merges the retrieved gestures by generating linking segments with rhythmic gestures. Different from [57], we observe that the gesture clip $G^{sem}$ retrieved from the semantic gesture repository is likely to have a different ongoing beat with the speech audio, especially in-the-wild ones. Thus, we prefer to inject the annotated ideal gesture frame rather than the entire segment, which enforces the generated rhythm to rely more on the actual input audio instead of $G^{sem}$. To identify a suitable injection timestamp for the ideal gesture frame, we propose an adaptive timestamp adjustment strategy based on the audio-gesture consistency score [28, 22] to guarantee our generated gesture process is synchronized with the semantic phrase, exempt from delay or rush. Specifically, we set the midpoint of $t_{s}$ and $t_{e}$ as the initial timestamp and repeatedly perform the RAG generation process while recording a temporary best score (see Section 4.5). We keep moving the timestamp following the binary search procedure until we find the one with the best score. According to the final best score, we achieve the most appropriate injection timestamp and a satisfactory generated result.
@@ 4.1 Datasets
In this section, we conducted comprehensive experiments on the SHOW [51] dataset and the proposed Streamer dataset to validate the effectiveness of our method.
SHOW Dataset.The SHOW dataset comprises 27 hours of talking videos from 4 actors, which includes reconstructed SMPL-X parameters of the whole body and synchronized speech data. The dataset emphasizes the synchronization between speech rhythm and gestures while lacking explicit semantic information. Following [29], we selected 10,706 clips as the training set and 1,313 clips as the test set.
@@ 4.2 Experimental Settings
Comparing Baselines. We select several state-of-the-art approaches, including TalkSHOW [51], Probtalk [29], and DiffuseStyleGesture (DSG) [49] as comparison methods. Both TalkShow and ProbTalk follow a VQ-VAE [39] paradigm, and TalkShow uses an autoregressive approach to predict motion distributions, while ProbTalk adopts a MaskGIT-like generative paradigm. DSG, on the other hand, is a diffusion-based approach. We use their pre-trained models for evaluation on the SHOW dataset and conduct training on the Streamer dataset using their released code. To ensure fairness, all comparisons are limited to the body and hands, excluding the face expressions.
Evaluation Metrics. We evaluate the performance of various methods from multiple perspectives. First, following [29], we select two commonly used metrics: Fréchet Gesture Distance (FGD) [52] and Beat Consistency Score (BC) [22]. Notably, we calculate the absolute difference for BC scores between predicted gestures and ground-truth (denoted as $\Delta BC$) for convenient comparison . Recent studies [9, 6] have pointed out that these metrics cannot fully and accurately reflect the quality of generated motions. Therefore, we propose two new evaluation metrics: Semantic Activation Rate (SAR) and Semantic Motion Distance (SMD) to assess the generation quality of semantic gestures from the Streamer dataset. SAR provides the success rate of generating correct gestures within the annotated semantic intervals, based on manual annotations. In terms of SMD, we first obtain the corresponding 3D key points from the generated gestures and the ground-truth gestures. Then we calculate the L1 distance and Dynamic Time Warping (DTW) distance on the upper limbs within each semantic segment, where DTW can handle motion sequences with varying rhythms. We randomly select 450 clips from each of the unseen and seen test sets to evaluate the SAR and SMD metrics. Notably, generated samples that are correctly activated but performed with the hand opposite to the ground truths, are filtered out when calculating SMD.

### v2

@@ Abstract
While increasing attention has been paid to co-speech gesture synthesis, most previous works neglect to investigate hand gestures with explicit and essential semantics. In this paper, we study co-speech gesture generation with an emphasis on specific hand gesture activation, which can deliver more instructional information than common body movements. To achieve this, we first build a high-quality dataset of 3D human body movements including a set of semantically explicit hand gestures that are commonly used by live streamers. Then we present a hybrid-modality gesture generation system GestureHYDRA built upon a hybrid-modality diffusion transformer architecture with novelly designed motion-style injective transformer layers, which enables advanced gesture modeling ability and versatile gesture operations. To guarantee these specific hand gestures can be activated, we introduce a cascaded retrieval-augmented generation strategy built upon a semantic gesture repository annotated for each subject and an adaptive audio-gesture synchronization mechanism, which substantially improves semantic gesture activation and production efficiency. Quantitative and qualitative experiments demonstrate that our proposed approach achieves superior performance over all the counterparts. The project page can be found at https://mumuwei.github.io/GestureHYDRA/.
@@ 1 Introduction
Co-speech gesture synthesis aims to produce gestures synchronized with speech audio, which enables a variety of applications in filmmaking, game design, robotics, and digital human creation [TALK-Act, showmaker, animateanyone, AudCast, remot]. Considerable enthusiasm [HA2G_ted, talkshow, diffstyle] is devoted to exploring the elusive relationship between speech audio and holistic body movements at the early stage.
Recently, a few studies [seeg, co-speech-3, wang2022vpu, beat, EMAGE, sege, SIGGesture] have shifted their focus on the semantic relevance within the gestures and pursue performance improvement by involving speech textual content as an additional condition. [beat, EMAGE] leverage word embedding from text scripts to perform semantic-relevant and rhythm-relevant co-speech motion modeling. The latest studies [sege, SIGGesture] prefer generative semantic retrieval via large language models (LLM) for better semantic injection. However, the semantic-gesture correlation studied in these approaches is not as clear as hand gestures with explicit instructions (e.g., numbers and directions), which can be frequently observed in our daily communication. Verifiably, hand gestures not only improve the efficiency of communication but also facilitate understanding via effective visual patterns [hands2minds, learn-to-count], which constructs a crucial component for co-speech generation. Therefore, to create life-like human behaviors that can efficiently deliver explicit information, a co-speech semantic gesture generation system is imperative.
@@ 2.1 Co-speech Gesture Generation
Human motion embodies rich and expressive information, rendering human motion analysis a longstanding and extensively studied research problem [gait, CLaM, EDGE, s3o]. Co-speech Gesture Generation focuses on understanding the complex relationship between speech audio and holistic body movements. Initially, researchers primarily relied on rule-based methods which required extensive manual effort to model this relationship [co-speech-rule1, co-speech-rule2, co-speech-rule3, co-speech-rule4, co-speech-rule5]. Data-driven methods [MDT-A2G, MambaTalk, SGToolkit, diffstyle, DiffSHEG, Audio2Gestures, DiffGesture, co-speech-Rhythmic, co-speech-2, co-speech-4-lstm, co-speech-5, DisCo] replaced rules with diverse models trained on large datasets, enabling automatic speech-gesture learning. VQ-VAE [vq-vae, pq-vae, tokengen] and diffusion models [ddim] have shown strong performance across downstream tasks [pixel2token, OOSTraj, dmdet] and are now the mainstream choices for co-speech gesture generation.
Some studies [talkshow, probtalk, diffstyle, DiffSHEG] focus on the rhythmic consistency between speech and generated body movements, particularly gestures, while roughly controlling limited motion styles using a one-shot identity vector. Recent studies [sege, SIGGesture, EMAGE] highlighted the importance of semantic gestures in conversational scenes and collected corresponding datasets. However, the semantic gestures in these works remain coarse and ambiguous. In contrast, our proposed large-scale Streamer dataset focuses on instructional and explicit semantic gestures (e.g., numbers and directions) in real-world scenarios, filling this gap.
@@ 2.2 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) is a hybrid approach that enhances generative models with external knowledge from databases. By retrieving relevant real-time data, RAG mitigates key limitations of LLMs [Llama3, Qwen25], such as knowledge cut-off and hallucination. Recent works [RAT, mR2AG] also apply RAG to text-to-image/video tasks by incorporating retrieved visual or textual context.
Recently, RAG has also been applied to gesture generation. ReMoDiffuse [ReMoDiffuse] explores the utilization of retrieved motions to provide useful guidance for the denoising process. For Co-speech generation, Semantic Gesticulator [sege] merges the retrieved gesture tokens into the initial token sequence via best merging time calculation. Similarly, SIGesture [SIGGesture] inserts the retrieved gestures into the denoised latent space for better generation. Unlike these methods that directly merge or insert the retrieval results back into the original gesture sequence, our approach relies on a cascaded RAG strategy, which first retrieves a key frame from relevant segments and re-generates expected rhythm and synchronized gestures with the input audio.
@@ 3.1 Task Formulation and Preliminaries
Task Formulation. Given an audio clip $A=\{a_{i}\}^{L-1}_{i=0}$, the objective of co-speech gesture generation is to reconstruct the target human gestures $G=\{g_{i}\}^{L-1}_{i=0}$ including natural body movements and responsive hand gestures with specific semantics, where $L$ denotes the sequence length. Concretely, we follow [diffstyle] to utilize a short clip with $S$ frames $G^{s}=\{g_{i}\}^{S-1}_{i=0}$ as the initial gestures (a.k.a., seed gestures).
Preliminaries for Diffusion models. Our system is built upon MDM[tevet2022humanmotiondiffusionmodel], which deploys diffusion transformer blocks for human motion modeling. The diffusion process operates as a Markov chain with $T$ steps, where noisy sequences $g_{1},...,g_{T}$ originate from real observations $g_{0}\sim q(g_{0})$. Gaussian noise is gradually applied at each step: $q\left(g_{t}\mid g_{t-1}\right):=\mathcal{N}\left(g_{t};\sqrt{1-\beta_{t}}g_{t-1},\beta_{t}\mathbf{I}\right)$. The reverse process aimed to reconstruct the original gestures from distorted gesture sequences. Unlike the image generation task [ldm], MDM prefers to predict the original sequence $g_{0}$ instead of the noise by using a learnable deep learning network, denoted by $\overline{g}_{0}=S_{\theta}\left(g_{t},t,c\right)$, where $c$ denotes the conditions. The training goal of this model is designed to minimize the mean squared error (MSE) between the original sequence $g_{0}$ and the predicted result $\overline{g}_{0}$, expressed as $L_{t}=\mathbb{E}_{g_{0},\epsilon,t}\left[g_{0}-S_{\theta}\left(g_{t},t,\mathrm{c}\right)\right]$.
@@ 3.2 Hybrid-Modality Diffusion Transformer
Hybrid Gesture Synthesis Training Procedure. During training, we first produce noisy gestures $G^{\prime}$ perturbed by Gaussian noise $\epsilon$ and leverage a Gesture Encoder $\bm{E}_{G}$ to extract noisy gesture latent $\mathbf{G}\in\mathbb{R}^{L\times D}$. Once the masking strategy is specified, we derive the key-frame gesture $G^{m}$ from $G$ by using the training mask $M=\{m_{i}\}^{L-1}_{i=0},m_{i}\in\{0,1\}$ and then perform channel-wise concatenation on $G^{m}$ and $M$. Then a KeyFrame Encoder $\bm{E}_{K}$ similar to $\bm{E}_{G}$ is employed to extract key-frame gesture latent, which is subsequently added to the noisy gesture latent $\mathbf{G}$ to form a fused latent $\mathbf{G^{K}}\in\mathbb{R}^{L\times D}$. In the next step, we turn our attention to the speech audio aspect. We first leverage a widely used Audio Encoder [wav2vec] $\bm{E}_{A}$ to extract audio embeddings $\mathbf{A}\in\mathbb{R}^{L\times C}$. Then we follow [DiffSHEG] to perform cross-modality alignment by using a simple gesture-audio fusion (GAF) module with a residual block instead of cross-attention layers [attention], which consists of a layer normalization (LN) operation and a multiple-layer perception (MLP). The fused latent $\mathbf{G}^{F}\in\mathbb{R}^{L\times C}$ can be defined as:
@@
Motion-Style Injective Transformer Layer. We initialize our transformer layer following [attention] and observe that generated gestures fail to preserve their original styles without further fine-tuning on individual data. Most previous approaches [diffstyle, probtalk] adopt one-hot embeddings for speaker identity guidance, however, they cannot provide generalized performance, especially when the number of identities keeps increasing.
Recently, a zero-shot talking face generation approach [resyncer] built upon speech-driven facial mesh animation proposes to improve its generalization ability by using an additional motion sequence as talking style reference. We take inspiration from it and propose to involve an additional gesture sequence and newly designed injective transformer layers for a more robust motion-style injection. Specifically, the motion-style injective transformer layer shown in Fig. 2(d) adopts a standard combination of a self-attention layer and a feed-forward network (FFN) layer with two style injection layers placed right after them, respectively. The style injection layer is designed to integrate the benefits of both dynamic motion components and static motion components, where the dynamic component denotes the motion-style embedding $\mathbf{S_{d}}\in\mathbb{R}^{L\times D}$ derived from a style encoder $\bm{E}_{S}$ representing captured motion-style information from external input, while the static component is an internal learnable motion memory-bank $\mathbf{S_{s}}\in\mathbb{R}^{L_{m}\times D}$ similar as  [vpgc, vprq], memorizing all the motion-styles within the training data. The final motion-style embedding $\mathbf{S}\in\mathbb{R}^{(L_{m}+L)\times D}$ is achieved by concatenating $\mathbf{S_{d}}$ and $\mathbf{S_{s}}$ along the temporal dimension and the style-injection attention can be formulated as:
@@
Dataset System FGD $\downarrow$ $\Delta BC$ $\downarrow$ SAR $\uparrow$ SMD-L1 $\downarrow$ SMD-DTW $\downarrow$
@@
Training Objectives. In addition to MSE loss $\mathcal{L}_{t}$ described in 3.1, we also incorporate the velocity loss [DiffSHEG] $L_{vec}$ and a 3D key point loss $L_{kp}$ to ensure similar human motion prediction to stable and accurate 3D positions, which provides effective guidance for downstream applications such as controllable video generation. Since we directly predict the original gestures $\overline{x}_{0}$ rather than the noise, it is convenient to calculate these auxiliary losses. The velocity loss and the 3D keypoint loss are both formulated using L1 distance. The velocity loss measures discrepancies in velocities, while the 3D keypoint loss quantifies the gap between predicted and ground-truth 3D keypoints. The total loss function is a weighted sum of the three losses:
@@
System FGD $\downarrow$ $\Delta BC$ $\downarrow$
@@ 3.3 Cascaded-Synchronized RAG
A recent study [sege] attempts to obtain injection timestamps by detecting audio beats corresponding to the stroke phase of gestures and merges the retrieved gestures by generating linking segments with rhythmic gestures. Different from [sege], we observe that the gesture clip $G^{sem}$ retrieved from the semantic gesture repository is likely to have a different ongoing beat with the speech audio, especially in-the-wild ones. Thus, we prefer to inject the annotated ideal gesture frame rather than the entire segment, which enforces the generated rhythm to rely more on the actual input audio instead of $G^{sem}$. To identify a suitable injection timestamp for the ideal gesture frame, we propose an adaptive timestamp adjustment strategy based on the audio-gesture consistency score [liu2022learning, learndance] to guarantee our generated gesture process is synchronized with the semantic phrase, exempt from delay or rush. Specifically, we set the midpoint of $t_{s}$ and $t_{e}$ as the initial timestamp and repeatedly perform the RAG generation process while recording a temporary best score (see Section 4.5). We keep moving the timestamp following the binary search procedure until we find the one with the best score. According to the final best score, we achieve the most appropriate injection timestamp and a satisfactory generated result.
@@ 4.1 Datasets
In this section, we conducted comprehensive experiments on the SHOW [talkshow] dataset and the proposed Streamer dataset to validate the effectiveness of our method.
SHOW Dataset.The SHOW dataset comprises 27 hours of talking videos from 4 actors, which includes reconstructed SMPL-X parameters of the whole body and synchronized speech data. The dataset emphasizes the synchronization between speech rhythm and gestures while lacking explicit semantic information. Following [probtalk], we selected 10,706 clips as the training set and 1,313 clips as the test set.
@@ 4.2 Experimental Settings
Comparing Baselines. We select several state-of-the-art approaches, including TalkSHOW [talkshow], Probtalk [probtalk], and DiffuseStyleGesture (DSG) [diffstyle] as comparison methods. Both TalkShow and ProbTalk follow a VQ-VAE [vq-vae] paradigm, and TalkShow uses an autoregressive approach to predict motion distributions, while ProbTalk adopts a MaskGIT-like generative paradigm. DSG, on the other hand, is a diffusion-based approach. We use their pre-trained models for evaluation on the SHOW dataset and conduct training on the Streamer dataset using their released code. To ensure fairness, all comparisons are limited to the body and hands, excluding the face expressions.
Evaluation Metrics. We evaluate the performance of various methods from multiple perspectives. First, following [probtalk], we select two commonly used metrics: Fréchet Gesture Distance (FGD) [co-speech-1] and Beat Consistency Score (BC) [learndance]. Notably, we calculate the absolute difference for BC scores between predicted gestures and ground-truth (denoted as $\Delta BC$) for convenient comparison . Recent studies [SIGGesture, DiffSHEG] have pointed out that these metrics cannot fully and accurately reflect the quality of generated motions. Therefore, we propose two new evaluation metrics: Semantic Activation Rate (SAR) and Semantic Motion Distance (SMD) to assess the generation quality of semantic gestures from the Streamer dataset. SAR provides the success rate of generating correct gestures within the annotated semantic intervals, based on manual annotations. In terms of SMD, we first obtain the corresponding 3D key points from the generated gestures and the ground-truth gestures. Then we calculate the L1 distance and Dynamic Time Warping (DTW) distance on the upper limbs within each semantic segment, where DTW can handle motion sequences with varying rhythms. We randomly select 450 clips from each of the unseen and seen test sets to evaluate the SAR and SMD metrics. Notably, generated samples that are correctly activated but performed with the hand opposite to the ground truths are filtered out when calculating SMD.